OpenStudy (anonymous):

simplify sqrt 27x^4/ sqrt 3x

OpenStudy (watchmath):

\(\sqrt{27x^4/(3x)}=\sqrt{9x^3}=3x\sqrt{x}\) if \(x\) is positive.